Direct Materials, Direct Labor, and Factory Overhead Cost Variance Analysis

Mackinaw Inc. processes a base chemical into plastic. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 78,000 units of product were as follows:

Standard Costs Actual Costs
Direct materials 218,400 lbs. at $4.80 216,200 lbs. at $4.60
Direct labor 19,500 hrs. at $18.10 19,950 hrs. at $18.40
Factory overhead Rates per direct labor hr.,
based on 100% of normal
capacity of 20,350 direct
labor hrs.:
Variable cost, $4.80 $92,660 variable cost
Fixed cost, $7.60 $154,660 fixed cost

Each unit requires 0.25 hour of direct labor.

Required:

a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Material Price Variance $
Direct Materials Quantity Variance $
Total Direct Materials Cost Variance $

b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Labor Rate Variance $
Direct Labor Time Variance $
Total Direct Labor Cost Variance $

c. Determine variable factory overhead controllable variance, the fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Variable factory overhead controllable variance $
Fixed factory overhead volume variance $
Total factory overhead cost variance $
Answer #1
Calculation of direct material Price variance
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($4.80/Pound - $4.60/Pound ) × 216200 Pound = $43240 F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
(218400-216200Pound )X $4.80 = $10560 F
Direct Mterial Cost Variance= Material Rate Avriance+ Material Quantity Variance
=$43240+10560=$53800
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($18.10/hour - $18.40/hour) × 19950 Hours= $5985 UF
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard Rate
= (19500 - 19950) × $18.10 = $8145 U F
Direct Labour Cost Variance= LRV+ LEV
=$5985UF+ $8145 UF=$14130 UF
Variable Overhead Controllable Variance= Std Cost Allowed - Actual Cost
( Actual Unit X std Hour X Std Rate)- Actual Cost Allowed
(78000**0.25*$4.80)-92660=$940 F
Fixed OH Volume Variance=(( Std hour for Actual Output)- Budgeted Hour)X Std Rate
((78000*0.25)-20350)X7.60=$6460 UF
Total Factory OH Cost variance= Variable Overhead Controllable variance+ FOH Volume Variance
=$940F+ $6460 UF=$5520 UF